OpenStudy (anonymous):

Complete the standard form of the equation -4/3y=3-x . x + y = 9

OpenStudy (nincompoop):

complete how?

OpenStudy (nincompoop):

do you want it in the form of Ax+by=c or y = mx+b

OpenStudy (anonymous):

what x and y

OpenStudy (nincompoop):

write the question properly and provide the equations properly

OpenStudy (anonymous):

that what i need help with

OpenStudy (nincompoop):

you have two equations, so that is basically solving for system of equation x and y

OpenStudy (anonymous):

4 and 3

OpenStudy (anonymous):

..

OpenStudy (nincompoop):

set up your equations properly into systems of equation \((1)~~\large \frac{4}{3}y=3-x \\ (2)~~ x+y = 9 \) what you want to do is multiply the entire first equation by \( \large \frac{3}{4} \) so it will look like \( \frac{3}{4} \times \frac{4}{3}y = \frac{3}{4} \times (3-x) \) alright go ahead and simplify the first equation

OpenStudy (anonymous):

3x+4y=9

OpenStudy (nincompoop):

what on earth are you doing? pay attention

OpenStudy (nincompoop):

I said simplify

OpenStudy (anonymous):

oh how u do that

OpenStudy (nincompoop):

do you not know how to multiply?

OpenStudy (anonymous):

ya u mean cross multiply

OpenStudy (nincompoop):

what does that even mean? cross multiply? cross multiply what?

OpenStudy (anonymous):

srry can u explain 2 me

OpenStudy (nincompoop):

if you can't simplify what I asked you to simplify, I think you have no business taking this algebra course. you should enroll into a pre-algebra or something lower.

OpenStudy (anonymous):

i dont get it

OpenStudy (nincompoop):

exactly, because you are not equipped to with the minimum knowledge to be in this course

OpenStudy (nincompoop):

good luck with your studies

OpenStudy (anonymous):

plz

OpenStudy (nincompoop):

buy yourself a $13 book called http://www.slideshare.net/santoshMrdexter/master-math-basic-math-and-pre-algebra-teach-yourself

OpenStudy (anonymous):

:L

OpenStudy (anonymous):

kirby help

OpenStudy (nincompoop):

or use online resources like ck.org http://www.ck12.org/book/Algebra-Explorations-Concepts-Pre-K-through-Grade-7/ and videos like khan academy